OpenStudy (anonymous):

MEDAL AND FAN PLEASE HELP! The radius of a cylinder is 3.5 ft. The height is 14 ft. Find the surface area and volume of the cylinder to the nearest tenth of a foot. Show your work please?

OpenStudy (anonymous):

surface area = 346.2 ft volume = 538.8 ft

OpenStudy (***[isuru]***):

the volume of a cylinder = area of the base ( circle) x heoght now the area of the circular base = pi x r^2 where r = radius pi = 3.14 ( approximately) = 3.14 x 3.5 x 3.15 = 38.465 sq.ft now to find the volume we have to multiply that value by the height volume = 38.465 x 14 = 538.51 cubic ft. if we round off it to the nearest tenth we will get volume = 538. 5 cubic ft. now to find the surface area Area = 2 x area of the circle + area of the curved part area of the curved part = circumference of the circle x height = 2 pi r x 14 = 2 x 3.14 x 3.5 x 14 sq. ft = 307 . 72 sq. ft now the total area = 307.72 + 38.465 sq ft = 346 . 185 if we round off it we will get 346.2 square ft. hope this will help ya :)

OpenStudy (anonymous):

Thank you :3

OpenStudy (***[isuru]***):

u r welcome:)!